# Neural network for excess noise estimation in continuous-variable quantum key distribution under composable finite-size security

### v2

@@ V.2 Network Architecture
The architecture of the network is illustrated in Fig.  1. The main goal here was to test the framework using a simple architecture, which comprises:
@@ V.3 Training Strategy
The network is trained using supervised learning, where the target values are derived from synthetic data generated via an additive white Gaussian noise (AWGN) model (see Eq. (1)). The predicted output $\hat{\mathrm{Y}}$ corresponds to the scaled noise term $a^{2}\hat{t}^{2}\hat{\xi}$, allowing a later inversion to recover the physical excess noise. The loss function is computed using the mean square error to ensure the minimization of Eq. (14). The model is optimized using the Adamw optimizer Loshchilov and Hutter (2019) via the Optax library, leveraging the JAX and Flax frameworks for high-performance computation.
@@ VI Numerical investigations
We investigate a finite-size security analysis employing neural networks within a specific architecture that leverages the signals $\{y_{i}\}_{m}$ and $\{x_{i}\}_{m}$ required for parameter estimation. A total of $10^{5}$ transmissivity values $t$ are sampled, each associated with a corresponding noise variance $\sigma^{2}=1+t^{2}\xi$. For every sampled pair, we generate computationally $N$ signal using the discussed protocol. Channel parameters are then estimated using $m\equiv N/2$ signals in the PEP 1 and PEP 2, which implement MLE and a neural network-based approach, respectively. This results in $10^{5}$ estimates for both $\sigma^{2}_{\text{max-MLE}}$ and $\sigma^{2}_{\text{max-NN}}$.
As an initial benchmark, we compare the precision of the estimators by analyzing the standard deviation between the predicted and true values of $\sigma^{2}$. As shown in Fig. 2, the neural network consistently yields lower deviations than the conventional MLE, reflecting the effectiveness of the error minimization performed during training. This increased accuracy stems from the fact that the network parameters are optimized to reduce a cost function, such as the mean squared error, which directly penalizes large prediction errors. Furthermore, one can verify that $\sigma^{2}_{\text{max}}\rightarrow\sigma^{2}$ as the sample size $m$ increases, as expected from asymptotic consistency of both estimators (see Eq. (10) and Eq. (24)).
@@
However, the main challenge is not simply to show that neural networks can be more precise (a result already discussed in literature Liu et al. (2018); Chin et al. (2021); Luo et al. (2022)), but demonstrate that they can also be $\epsilon_{PE}\text{-secure}$ for parameter estimation in CV-QKD. An estimate is $\epsilon_{PE}\text{-secure}$ if, and only if, all the points estimated are inside the confidence intervals with probability at least $1-\epsilon_{PE}/2$, i.e., one can never estimate $\sigma^{2}_{\text{max}}$ below its real value considering this probability. Figure 3 illustrates this behavior by depicting the average trend line obtained from the closest predicted points to the ideal reference, computed across all samples. This visualization emphasizes the proximity between predicted and true values. Notably, in all cases, the estimates remained above the ideal curve, ensuring that no underestimation occurred throughout the tested configurations.
@@
The precision of parameter estimation plays a critical role in the secret-key rate — since we must overestimate the excess noise with high confidence, more accurate estimators yield smaller values of $\xi$, ensuring that $\Gamma_{\epsilon_{PE}}\in\mathcal{C}_{\epsilon_{PE}}$. The results in Fig. 4 illustrate this behavior: the estimated secret-key rate consistently remains below the ideal rate, as expected. This outcome confirms the operational security of the parameter estimation procedure described in PEP 2, with the advantage of achieving higher rates.
Figure 4: Secret-key rate using the discussed protocol with parameters described in Tab. LABEL:tab:parameters. Dot-dashed line, dashed line and dotted line corresponds, respectively, to $N=2\cdot 10^{6}$, $2\cdot 10^{7}$ and $2\cdot 10^{8}$ signals. The use of neural networks allowed a gain of $6.1$ km, $1.3$ km and $3.0$ km, respectively. In all cases, the estimated values was never superior to the real values
@@ VII Conclusion
Neural networks have been increasingly adopted as estimation tools in quantum information. In the context of QKD, the trade-off between computational cost and estimation accuracy becomes particularly relevant. While neural networks typically demand greater computational resources compared to conventional methods, they can offer improved precision. This is significant in the finite-size regime, where even modest gains in parameter estimation may translate into a positive key rate in otherwise insecure regimes.
In conclusion, this article gives a finite-size analysis for secure CV-QKD using networks for excess noise estimation. While the neural network employed in our simulations demonstrates improved estimation accuracy compared to the conventional MLE, we emphasize that the primary contribution of this work lies not in the superiority of a specific architecture, but in demonstrating that neural network-based estimators can be incorporated into parameter estimation routines for CV-QKD without compromising composable security. Although more robust or efficient architectures may be developed, our findings indicate that such data-driven approaches are compatible with finite-size security proofs. This insight enables the use of flexible and potentially adaptive estimation strategies in practical QKD systems, paving the way for further integration of machine learning techniques into secure quantum communication protocols.
@@ Acknowledgements.
This work has been fully funded by the project Computational Architecture for Flexible QKD System Post-processing Platform supported by QuIIN - Quantum Industrial Innovation, EMBRAPII CIMATEC Competence Center in Quantum Technologies, with financial resources from the PPI IoT/Manufatura 4.0 of the MCTI grant number 053/2023, signed with EMBRAPII.
@@ Appendix A Computational cost of the neural network
The computational complexity of the neural architecture is primarily determined by the number of trainable parameters and the per-sample inference cost. The model is implemented in Flax and trained using the JAX framework, leveraging XLA compilation and hardware acceleration for efficient execution. Let $n$ be the number of samples per iteration and $d$ the input dimension, with $d=6$ corresponding to the MLE estimate, mean values, variances, and covariances extracted from the AWGN model. The network processes input tensors of shape $(n,d)$.

### v3

@@ V.2 Network Architecture
The architecture of the network is illustrated in Fig. 2. The main goal here was to test the framework using a simple architecture, which comprises:
@@ V.3 Training Strategy
The neural network is trained using supervised learning, where the target values are derived from synthetic data generated via a gaussian channel model (see Eq. (1)). At each step of the training loop, a new batch of data is produced by drawing the channel parameters uniformly at random: the transmittance $t\in[0.01,1]$ and the excess noise $\xi\in[0.0,0.1]$ SNU. Specifically, 64 distinct transmittance values are drawn per loss evaluation, each paired with a random excess noise value, and this process repeats until convergence. This procedure ensures that the network is exposed to a continuously refreshed and representative sample of the channel parameter space, promoting robust generalization without overfitting to a fixed dataset.
The loss function is defined as the mean squared error, ensuring the minimization of Eq. (16) and consistency with the statistical framework introduced in Sec. IV. The model parameters are optimized using the AdamW optimizer Loshchilov and Hutter (2019) implemented via the Optax library, leveraging the JAX and Flax frameworks for high-performance computation. Further details regarding the optimization procedure and hyperparameter settings are provided in Appendix D.
@@ VI Numerical investigations
We investigate a finite-size security analysis employing neural networks within a specific architecture that leverages the signals $\{y_{i}\}_{m}$ and $\{x_{i}\}_{m}$ required for parameter estimation. A total of $10^{5}$ transmittance values $t$ are sampled, each associated with a corresponding noise variance $\sigma^{2}=1+t^{2}\xi$. For every sampled pair, we generate computationally $N$ signal using the discussed protocol. Channel parameters are then estimated using $m\equiv N/2$ signals in the PEP III and PEP IV, which implement MLE and a neural network-based approach, respectively. This results in $10^{5}$ estimates for both $\sigma^{2}_{\text{max-MLE}}$ and $\sigma^{2}_{\text{max-NN}}$.
The synthetic data used in our numerical investigations are generated according to the AWGN model (Eq. (1)). This approach is standard for finite-size security analysis, as the gaussian attack represents the worst-case scenario for CV-QKD under the assumed channel model Wolf et al. (2006); Jain et al. (2022). By using synthetic data, we maintain full control over channel parameters, enabling a precise assessment of the neural-network estimator’s statistical performance under the most adversarial conditions. While experimental data could introduce additional non‑Gaussian imperfections, the synthetic‑data approach provides a fundamental validation of the estimator’s statistical properties and its compliance with finite‑size security proofs.
As an initial benchmark, we compare the precision of the estimators by analyzing the standard deviation between the predicted and true values of $\sigma^{2}$. As shown in Fig. 3, the neural network consistently yields lower deviations than the conventional MLE, reflecting the effectiveness of the error minimization performed during training. This increased accuracy stems from the fact that the network parameters are optimized to reduce a cost function (Eq. (16)), which directly penalizes large prediction errors. Furthermore, one can verify that $\sigma^{2}_{\text{max}}\rightarrow\sigma^{2}$ as the sample size $m$ increases, as expected from asymptotic consistency of both estimators (see Eq. (10) and Eq. (26)).
@@
However, the main challenge is not simply to show that neural networks can be more precise (a result already discussed in literature Liu et al. (2018); Chin et al. (2021); Luo et al. (2022)), but demonstrate that they can also be $\epsilon_{PE}\text{-secure}$ for parameter estimation in CV-QKD. An estimate is $\epsilon_{PE}\text{-secure}$ if, and only if, all the points estimated are inside the confidence intervals with probability at least $1-\epsilon_{PE}/2$, i.e., one can never estimate $\sigma^{2}_{\text{max}}$ below its real value considering this probability. Figure 4 illustrates this behavior by depicting the average trend line obtained from the closest predicted points to the ideal reference, computed across all samples. The results empirically validates this operational equivalence, demonstrating that, for all tested sample sizes, the neural estimate $\sigma^{2}_{\text{max-NN}}$ never falls below the true value $\sigma^{2}$, thereby satisfying the security condition with the specified probability $\epsilon_{PE}$.
@@
The precision of parameter estimation plays a critical role in the secret-key rate — since we must overestimate the excess noise with high confidence, more accurate estimators yield smaller values of $\xi$, ensuring that $\Gamma_{\epsilon_{PE}}\in\mathcal{C}_{\epsilon_{PE}}$. The results in Fig. 5 illustrate this behavior: the estimated secret-key rate consistently remains below the ideal rate, as expected. This outcome confirms the operational security of the parameter estimation procedure described in PEP IV, with the advantage of achieving higher rates.
Figure 5: Secret-key rate using the discussed protocol with parameters described in Tab. LABEL:tab:parameters. Dot-dashed line, dashed line and dotted line corresponds, respectively, to $N=2\cdot 10^{6}$, $2\cdot 10^{7}$ and $2\cdot 10^{8}$ signals. The use of neural networks allowed a gain of $6.1$ km, $1.3$ km and $3.0$ km, respectively. In all cases, the estimated values was never superior to the real values.
@@ VII Conclusion
Neural networks are increasingly being adopted as estimation tools in quantum information, with their application in QKD highlighting a critical balance between computational cost and estimation accuracy. While these models typically require greater resources than conventional methods, they can achieve substantially improved precision. This enhancement is particularly significant in the finite size regime, where even modest gains in parameter estimation can enable a positive secret key rate in otherwise insecure operational regimes. The neural network estimator presented here demonstrates how machine learning can be securely integrated to enhance practical QKD performance. By learning subtle statistical dependencies within correlated quadrature data, the network delivers tighter bounds on excess noise than standard linear estimators, directly leading to higher secret-key rates and extended transmission distances. This improvement is especially valuable when the number of transmitted signals is limited, as it can reduce the parameter estimation overhead and extend the secure range of existing systems.
The developed framework is broadly applicable to Gaussian modulated CV QKD under collective attacks and paves the way for future adaptations, including discrete modulation, time varying channels, or hybrid schemes incorporating real time feedback for adaptive estimation. Considering our approach, the neural network estimator provides tighter parameter bounds for collective Gaussian attacks, which form the basis for security against coherent attacks via reduction techniques. Since these reductions rely only on the validity of the worst-case bounds, our method can be seamlessly integrated into composable security proofs against general attacks. Thus, the gains in key rate and distance demonstrated in this work are preserved while maintaining the same security level against coherent attacks.
In conclusion, this article provides a finite-size analysis for secure CV-QKD using networks for excess noise estimation. While the neural network employed in our simulations demonstrates improved estimation accuracy compared to the conventional MLE, we emphasize that the primary contribution of this work lies not in the superiority of a specific architecture, but in demonstrating that neural network-based estimators can be incorporated into parameter estimation routines for CV-QKD without compromising composable security. Although more robust or efficient architectures may be developed, our findings indicate that such data-driven approaches are compatible with finite-size security proofs. This insight enables the use of flexible and potentially adaptive estimation strategies in practical QKD systems, paving the way for further integration of machine learning techniques into secure quantum communication protocols.
@@ Acknowledgements.
This work has been fully funded by the project Computational Architecture for Flexible QKD System Post-processing Platform and QRNG: Post-processing and Experimental supported by QuIIN - Quantum Industrial Innovation, EMBRAPII CIMATEC Competence Center in Quantum Technologies, with financial resources from the PPI IoT/Manufatura 4.0 of the MCTI grant number 053/2023, signed with EMBRAPII. MD thanks financing from the European Union (HORIZON-MSCA-2023 Postdoctoral Fellowship, 101153602 - COCoVaQ).
@@ Appendix A Computational cost of the neural network
The computational complexity of the neural architecture is primarily determined by the number of trainable parameters and the per-sample inference cost. The model is implemented in Flax and trained using the JAX framework, leveraging XLA compilation and hardware acceleration for efficient execution. Let $n$ be the number of samples per iteration and $d$ the input dimension, with $d={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}6}$ corresponding to the MLE estimate, mean values, variances, and covariances extracted from the AWGN model. The network processes input tensors of shape $(n,d)$.
@@
Appendix B Practical implementation for confidence interval computation
[…]
Regularization of the matrix inversion. In practice, the matrix $\mathbf{F}^{T}\mathbf{F}$ is often ill-conditioned, especially when the number of samples $m$ is not substantially larger than the number of parameters $p$. The observed ill‑conditioning arises because a significant proportion of its eigenvalues lie extremely close to zero. In practice, finite‑precision arithmetic can produce a small number of eigenvalues with a negligible negative magnitude ($\sim 10^{-10}$), which we treat as effectively zero. The presence of these near‑zero eigenvalues renders the matrix nearly singular and causes its inversion to be numerically unstable.
To ensure numerical stability when computing the inverse $(\mathbf{F}^{T}\mathbf{F})^{-1}$, we employ Tikhonov regularization. Specifically, we replace $(\mathbf{F}^{T}\mathbf{F})^{-1}$ with $(\mathbf{F}^{T}\mathbf{F}+\lambda\mathbf{I})^{-1}$, where $\lambda>0$ is a small regularization parameter and $\mathbf{I}$ is the $p\times p$ identity matrix. For all simulations reported in this work, we set $\lambda=10^{-4}$ , resulting in a condition number $\kappa\approx 4.5\cdot 10^{3}$. This value was chosen via cross-validation on a separate validation set, aiming to minimize the mean squared error of the variance predictions while maintaining well-conditioned matrix inversions. Therefore, the regularization transforms the original ill-conditioned matrix $\mathbf{F}^{T}\mathbf{F}$ (with $\sim 15\%$ of its eigenvalues below $\lambda$) into the regularized matrix $\mathbf{F}^{T}\mathbf{F}+\lambda\mathbf{I}$, which has all eigenvalues $\geq\lambda$, eliminating any numerical instability and practical rank deficiency arising from near-zero eigenvalues.
Appendix C Inference latency and real‑time feasibility
To generate Mbit/s of security keys, a practical real‑time CV‑QKD system should be able to process tens or hundreds of MBaud Zhang et al. (2024). In this context, parameter estimation must be completed within stringent real‑time limits, typically processing a portion of that system throughput. The neural network estimator introduced in this work would be able to meets this requirement by operating exclusively in inference mode once offline training is completed.
A single forward pass determines its computational demand through a fully‑connected network with $p=4450$ parameters, which translates to $8679$ elementary arithmetic operations per frame. This computational cost is primarily driven by matrix‑vector multiplications in the hidden layers. It is constant and does not scale with the number of symbols processed per frame, being negligible compared to the required data preprocessing to calculate its inputs (27), which include the statistical moments and the transmittance MLE (8).
Consequently, the computational burden is dominated by the linear complexity $\mathcal{O}(m)$ required to prepare the statistical inputs. Both the conventional analytic method and the neural-network approach must first compute $\hat{t}_{MLE}$, a step whose cost is common to both. The divergence occurs in the estimation of the variance and excess noise. The classical method computes the sum of squared residuals, iterating once over the data with the already obtained $\hat{t}_{MLE}$. The neural-network method, in turn, requires the calculation of five sufficient statistics (means, variances, and a covariance) from the raw data vectors. As detailed in Table 2, the cost of generating these statistical inputs for a frame of $m=10^{8}$ symbols is linear and comparable between the two methods. The key distinction emerges after this stage: while the analytic method proceeds with a final arithmetic step, the neural-network approach performs a single forward pass through a fixed architecture. This inference requires only $\approx 8.7\times 10^{3}$ operations, which is a negligible, constant overhead that is effectively imperceptible against the $\mathcal{O}(10^{8})$ operations already expended on the shared transmittance estimation and statistical aggregation.
Table 2: Computational cost per frame for parameter estimation with $m=10^{8}$ correlated measurement pairs. The table compares the dominant, shared cost of transmittance estimation with the subsequent steps for noise variance estimation, highlighting the fixed and negligible overhead introduced by the neural-network inference. All costs are estimated in terms of arithmetic operations.
Processing Stage & Scope Est. Cost (Ops)
Transmittance Estimation (Shared Input)
Calculation of $\hat{t}_{MLE}$ $\mathcal{O}(m)\approx 2\times 10^{8}$
   Divisions* $1\times 10^{8}$
Variance Estimation
Analytic (MLE) Approach:
   Residual computation $\sum(y_{i}-\hat{t}x_{i})^{2}$ $\mathcal{O}(m)\approx 4.0\times 10^{8}$
Neural-Network-Based Approach:
   Preparation of 5 sufficient statistics $\mathcal{O}(m)\approx 8.0\times 10^{8}$
   Neural-network inference (single forward) $\mathbf{\mathcal{O}(1)\approx 8.7\times 10^{3}}$
    Division operations are computationally expensive and can take many cycles depending on the hardware support.
In terms of execution time, a single forward pass executes well below one millisecond on a standard CPU and can be accelerated to microsecond latency when deployed on embedded processors, FPGAs, or dedicated digital signal processing platforms. In a test performed on an NVIDIA RTX3090 GPU, with $10000$ runs, the average inference time was $329.65\mu s$, with a standard deviation of $40.20\mu s$.
Notably, no online training or backpropagation is performed during operation. Furthermore, the Jacobian matrix required for constructing the worst‑case confidence interval is pre‑computed offline using a large synthetic dataset and reused at runtime. These characteristics confirm that the neural‑network estimator is not only statistically sound but also practically viable for integration into real‑world CV‑QKD systems, where low‑latency and deterministic processing are essential.
Appendix D Optimization details
The network parameters were optimized using the AdamW optimizer Loshchilov and Hutter (2019), implemented via the Optax library within the JAX and Flax frameworks. An exponential learning-rate decay schedule was employed, with an initial learning rate of $1\times 10^{-4}$, transition steps of 500, and a decay rate of 0.9. The weight decay coefficient was set to $1\times 10^{-5}$.
Training was performed using a batch size of 64 samples for 15 epochs, with 512 gradient-update steps per epoch. For numerical stability, gradients were clipped by global norm with a threshold of 0.5. Model performance was periodically evaluated every 256 training steps. No adaptive hyperparameter tuning, learning-rate warm-up, or early stopping strategy was applied. All experiments were conducted using a fixed JAX pseudo-random number generator seed to ensure reproducibility.